an accountant can pick 3 out of 50 tax returns to get a special audit. how many ways can these 3 be picked if the order does matter?
Answer by stanbon(75887)

an accountant can pick 3 out of 50 tax returns to get a special audit. how many ways can these 3 be picked if the order does matter?
-----
Ans: 50P3 = 50!/(50-3)! = 50!/47! = 50*49*48 = 117,600
=============================
Cheers,
Stan H.
=============================